# Towards Foundation Model for Chemical Reactor Modeling: Meta-Learning with Physics-Informed Adaptation

### v2

@@ 5 Few-Shot Performance to Fixed-Order Unseen Reactions
We initiate the meta-training phase by training the Reptile-based foundation model on 500 CSTRs, 500 BRs, and 500 PFRs with different parameters representing various reactions in different settings of reactors (e.g., reactor size, flow rate, etc.). In actual deployment to industrial engineering processes, the foundation model is trained offline, and the costs associated with the meta-training process, such as dataset generation, are not a primary concern. Following this, in the meta-testing phase, we evaluate the few-shot performance of the pre-trained foundation model on 5 previously unseen CSTRs, 5 previously unseen BRs, and 5 previously unseen PFRs, respectively, using MSE as the evaluation metric. Specifically, during meta-testing, we fine-tune the foundation model using a limited number of samples from the designated reaction and assess its performance by computing the MSE over all samples from the same reaction, where the samples used for few-shot fine-tuning and the collocation points are randomly selected from the whole operational region of the designated reaction. In Figs. 3, 4, 5, and 7, the solid line represents the mean of 5 trials on different unseen reactions, while the shaded region indicates the standard deviation.
We employ a 2-layer RNN with 64 neurons per layer and ReLU activation functions as the foundation model architecture. The model is optimized using the Adam optimizer. In the following subsections, we will provide a comprehensive analysis of the few-shot performance with respect to five different approaches on unseen CSTRs, BRs, and PFRs respectively, formulated as follows:
@@
    Reptile-based foundation model with normal data-driven adaptation: In this method, the model is first pre-trained on the 500 CSTRs, 500 BRs, and 500 PFRs dataset using the Reptile learning rule. After pre-training, the foundation model is fine-tuned with few-shot samples in a normal data-driven manner.
@@
    Reptile-based foundation model with physics-informed adaptation: This method combines the benefits of both Reptile-based meta-learning and physics-informed learning. The model is initially pre-trained on the 500 CSTRs, 500 BRs, and 500 PFRs dataset using the Reptile learning rule to enhance its adaptability. Then, during fine-tuning with few-shot samples, a physics-informed loss function is employed to leverage domain-specific knowledge.
These approaches vary in their use of prior knowledge, meta-learning techniques, and incorporation of physics-based constraints. By comparing their few-shot performance, we can assess which method is most effective for the given tasks and dataset. The experiments were conducted on Intel Core i7-12700 processor with 64 GB of RAM and a single run for meta-training took about 48 hours, using Python and TensorFlow Keras (Chollet et al., 2015). Moreover, the source code is available in https://github.com/killingbear999/chemical-reactor-foundation-model.
@@ 5.1 Few-Shot Adaptation to Unseen CSTRs
Next, we explore how the number of shots influences model performance across the five aforementioned approaches (i.e., we use 100 collocation points for PI-based methods). From Fig. 3(b), it is evident that the transfer learning approach exhibits the poorest performance. This is because the model encounters confusion during the pre-training phase using conventional learning algorithms, given the disparate distributions of CSTRs, BRs, and PFRs. Conventional physics-informed learning (i.e., without Reptile model) starts with a strong foundation but halts its learning process due to the limited additional information provided by the constrained number of shots, eventually matching the performance of normal data-driven learning with around 50 shots. Reptile with normal adaptation surpasses physics-informed learning after around 30 shots and requires more than 15 shots to achieve satisfactory performance (i.e., reaching an MSE magnitude of $10^{-3}$). In contrast, Reptile with physics-informed adaptation reaches a comparable level with just about 1 shot. Notably, Reptile with physics-informed adaptation maintains the lowest MSE consistently as the number of shots increases.
@@ 5.2 Few-Shot Adaptation to Unseen BRs
Next, we investigate how the number of shots influences model performance across the five aforementioned approaches (i.e., using 100 collocation points for PI-based methods). From Fig. 4(b), it is apparent that the transfer learning approach performs the poorest, due to confusion during the pre-training phase caused by the distinct distributions of CSTRs, BRs, and PFRs. Conventional physics-informed learning begins with a robust foundation but stalls in its learning trajectory due to the limited additional information from the constrained number of shots, eventually matching normal data-driven learning performance with approximately 25 shots. More importantly, consistent with earlier findings, it is more challenging to train BRs using PI-based methods compared to CSTRs. Reptile with normal adaptation requires approximately 40 shots to achieve satisfactory performance (i.e., reaching an MSE magnitude of $10^{-3}$). Conversely, Reptile with physics-informed adaptation attains a similar level with only about 25 shots. Notably, Reptile with physics-informed adaptation consistently maintains the lowest MSE as the number of shots increases.
@@ 5.3 Few-Shot Adaptation to Unseen PFRs
Comparing Fig. 3(b), Fig. 4(b), and Fig. 5(b), we can observe that the foundation model prioritizes reaction types in the following order: CSTRs $>$ PFRs $>$ BRs. As a result, during the adaptation phase, CSTRs achieve the best performance, followed by PFRs, and then BRs. The performance of the different methods is summarized as follows: Reptile with physics-informed adaptation $>$ Reptile with normal adaptation $>$ Physics-informed learning $>$ Normal data-driven learning $>$ Transfer learning.
@@ 6 Extension to Varying Integer Orders of Reactions
Next, we extend the proposed method to handle varying integer reaction orders. Specifically, we meta-train foundation models for different integer orders of reactions. Consequently, for an unseen integer order of reaction, we integrate their capabilities through ensemble learning, referred to as the ensemble model, with just a few data samples, without requiring precise knowledge of the reaction order, the reaction type, or the complete material and energy balance equations. The high-level system architecture of the ensemble model for varying integer reaction orders is illustrated in Fig. 6. It can be described in three steps: 1) Train $N$ separate foundation models, each corresponding to a different integer reaction order; 2) For a new reaction with an integer reaction order, perform few-shot adaptation on all $N$ foundation models to obtain $N$ intermediate models; 3) Apply min-voting (i.e., select the model with the lowest validation/testing MSE) to determine the final ensemble model, utilizing ensemble learning.
@@
    Incorrect order Reptile with normal adaptation: In this approach, after obtaining the foundation model via Reptile, an incorrect order of the foundation model (i.e., the reaction order used during meta-training differs from that of the unseen meta-testing data) is applied during the meta-testing phase using a standard data-driven adaptation method.
@@
    Incorrect order Reptile with physics-informed adaptation: Similar to the previous approach, an incorrect order of the foundation model is used during the meta-testing phase, but this time adaptation is performed using a physics-informed loss function.
@@
Notably, for CSTRs, the ensemble Reptile with normal adaptation performs worse than incorrect-order Reptile with physics-informed adaptation. This is likely because the foundation model is better suited to CSTRs than to PFRs and BRs, as observed in Fig. 3(b), Fig. 4(b), and Fig. 5(b). Thus, even with an incorrect reaction order in the foundation model, it can still quickly adapt to unseen CSTR reactions. In summary, for a well-trained foundation model (particularly with respect to CSTRs), incorporating a physics-informed loss term accelerates adaptation. However, for a foundation model that is not as well trained (e.g., with respect to BRs and PFRs), the physics-informed loss term can potentially decelerate adaptation, leading to increased confusion for the model.
@@
To further extend the proposed foundation model to incorporate varying fractional reaction orders, the meta-training dataset can be expanded to include random fractional reaction orders. During the meta-testing phase, physics-informed adaptation is then applied to facilitate accurate learning and adaptation.
@@ 7 Limitations and Future Works
The efficacy of our proposed method is heavily influenced by the challenges associated with training PI-based methods for specific tasks due to the use of estimated parameter values in the physics-informed loss terms. Notably, BRs and PFRs pose greater difficulty for PI-based training, resulting in less discernible differences between our Reptile-based foundation model with physics-informed adaptation and the one with normal data-driven adaptation, compared to CSTRs, where the distinction is more pronounced. To address this, inspired by numerical integration, we propose integrating the driving equations directly into the recurrent cells as a different way to incorporate physical prior into the model, rather than incorporating them into the loss term as a soft constraint. Another challenge with PI-based methods, as discussed in Remark 9, lies in the selection of collocation points.
Moreover, this work is simulation-based due to the lack of public datasets. Unlike computer vision and natural language processing, which benefit from publicly shared data through academic research and open-access journals, chemical engineering deals with large-scale industrial processes, where data collection is complex, proprietary, and highly variable. In future work, Aspen HYSYS (i.e., an industry-level chemical process simulator that models everything from unit operations to full chemical plants and refineries) can be used for data generation to further improve our foundation model by accounting for more training data from Aspen high-fidelity models. Additionally, experimental data from a real-world photochemical reactor (i.e., an unseen reactor type) will be used to assess whether the developed foundation model can adapt to it quickly. A successful demonstration could further extend the model’s applicability to unseen reactor types with unknown reaction orders.
@@ 8 Conclusion
In this study, we introduce a novel application of foundation models in nonlinear chemical process modeling for chemical reactors. Using meta-learning and physics-informed adaptation, we have developed a foundation model for rapid adaptation to previously unseen chemical reactions in CSTRs, BRs, and PFRs, with just a few training shots. The proposed method surpasses conventional approaches such as data-driven learning, physics-informed learning, transfer learning, and meta-learning in a few-shot setting. The successful implementation of the proposed foundation model holds the potential to revolutionize chemical reactor modeling by drastically reducing data requirements and improving generalizability to new reactions. This advancement could have significant implications for various applications in chemical engineering, such as process optimization and control.
@@ 9 Acknowledgments
Financial support from the A*STAR MTC YIRG 2022 Grant ( [^2] K3024) and NRF-CRP Grant 27-2021-0001 is gratefully acknowledged.

### v3

@@ 5 Few-Shot Performance to Fixed-Order Unseen Reactions
We initiate the meta-training phase by training the Reptile-based “foundation model” on 500 CSTRs, 500 BRs, and 500 PFRs with different parameters representing various reactions in different settings of reactors (e.g., reactor size, flow rate, etc.). In actual deployment to industrial engineering processes, the “foundation model” is trained offline, and the costs associated with the meta-training process, such as dataset generation, are not a primary concern. Following this, in the meta-testing phase, we evaluate the few-shot performance of the pre-trained “foundation model” on 5 previously unseen CSTRs, 5 previously unseen BRs, and 5 previously unseen PFRs, respectively, using MSE as the evaluation metric. Specifically, during meta-testing, we fine-tune the “foundation model” using a limited number of samples [^3] Each sample corresponds to a concentration-temperature pair at a single point in time, serving as an initial condition (i.e., the current state), while the RNN predicts the trajectory of future states. from the designated reaction and assess its performance by computing the MSE over all samples from the same reaction, where the samples used for few-shot fine-tuning and the collocation points are randomly selected from the whole operational region of the designated reaction. In Figs. 6, 7, 8, and 11, the solid line represents the mean of 5 trials on different unseen reactions, while the shaded region indicates the standard deviation. The $y$-axis is presented on a logarithmic scale for clarity.
We employ a 2-layer RNN with 64 neurons per layer and Tanh activation functions as the “foundation model” architecture. A simple RNN cell follows:
$h_{t}$ $=g_{1}(W^{(x)}x_{t}+\theta h_{t-1}+b^{(h)})$ (15a)
$y_{t}$ $=g_{2}(W^{(y)}h_{t}+b^{(y)})$ (15b)
where $h_{t}$ is the hidden state at time step $t$, $x_{t}$ is the input at time step $t$, $h_{t-1}$ is the hidden state from the previous time step, $W^{(x)}$, $\theta$ and $W^{(y)}$ are weight matrices for the input, hidden state, and output respectively, $b^{(h)}$ and $b^{(y)}$ are the bias vectors for the hidden state and output respectively, and $y_{t}$ is the output at time step $t$. The architecture of a 2-layer RNN is illustrated in Fig. 4. During meta-training, the following input-output configurations are utilized based on the type of training sample, summarized as follows: (1) for a CSTR, the inputs consist of $[T_{t},C_{A_{t}},\Delta Q_{t},\Delta C_{{A0}_{t}}]$ at the current time step $t$, and the outputs entail the state trajectory $[T_{t+1,\ldots,n},C_{A_{t+1,\ldots,n}}]$ over the subsequent $n$ time steps; (2) for a BR, the inputs consist of $[T_{t},C_{A_{t}},\Delta Q_{t},\Delta Q_{t}]$ at the current time step $t$, and the outputs entail the state trajectory $[T_{t+1,\ldots,n},C_{A_{t+1,\ldots,n}}]$ over the subsequent $n$ time steps; (3) for a PFR, the inputs consist of $[T_{t},C_{A_{t}},\Delta T_{c_{t}},\Delta T_{c_{t}}]$ at the current time step $t$, and the outputs entail the state trajectory $[T_{t+1,\ldots,n},C_{A_{t+1,\ldots,n}}]$ over the subsequent $n$ time steps.
Figure 4: A recurrent neural network and its unfolded structure.
The model is optimized using the Adam optimizer. In the following subsections, we will provide a comprehensive analysis of the few-shot performance with respect to five different approaches on unseen CSTRs, BRs, and PFRs respectively (see Fig. 5), formulated as follows:
@@
    Reptile-based “foundation model” with normal data-driven adaptation: In this method, the model is first pre-trained on the 500 CSTRs, 500 BRs, and 500 PFRs dataset using the Reptile learning rule. After pre-training, the “foundation model” is fine-tuned with few-shot samples in a normal data-driven manner.
@@
    Reptile-based “foundation model” with physics-informed adaptation: This method combines the benefits of both Reptile-based meta-learning and physics-informed learning. The model is initially pre-trained on the 500 CSTRs, 500 BRs, and 500 PFRs dataset using the Reptile learning rule to enhance its adaptability. Then, during fine-tuning with few-shot samples, a physics-informed loss function is employed to leverage domain-specific knowledge.
Figure 5: Summary of different approaches.
These approaches vary in their use of prior knowledge, meta-learning techniques, and incorporation of physics-based constraints. By comparing their few-shot performance, we can assess which method is most effective for the given tasks and dataset. The experiments were conducted on Intel Core i7-12700 processor with 64 GB of RAM and a single run for meta-training took about 48 hours, using Python and TensorFlow Keras. Moreover, the source code is available in https://github.com/killingbear999/chemical-reactor-foundation-model.
@@ 5.1 Few-Shot Adaptation to Unseen CSTRs
Next, we explore how the number of shots influences model performance across the five aforementioned approaches (i.e., we use 100 collocation points for PI-based methods). From Fig. 6(b), it is evident that the transfer learning approach exhibits the poorest performance. This is because during the pre-training phase, the model encounters difficulties in learning a unified representation due to the significantly different distributions of CSTRs, BRs, and PFRs when using conventional learning algorithms. Conventional physics-informed learning (i.e., without Reptile model) starts with a strong foundation but halts its learning process due to the limited additional information provided by the constrained number of shots, eventually matching the performance of normal data-driven learning with around 50 shots. Reptile with normal adaptation surpasses physics-informed learning after around 30 shots and requires more than 15 shots to achieve satisfactory performance (i.e., reaching an MSE magnitude of $10^{-3}$). In contrast, Reptile with physics-informed adaptation reaches a comparable level with just about 1 shot. Notably, Reptile with physics-informed adaptation maintains the lowest MSE consistently as the number of shots increases.
@@ 5.2 Few-Shot Adaptation to Unseen BRs
Next, we investigate how the number of shots influences model performance across the five aforementioned approaches (i.e., using 100 collocation points for PI-based methods). From Fig. 7(b), it is apparent that the transfer learning approach performs the poorest, due to difficulties in learning a unified representation during the pre-training phase caused by the distinct distributions of CSTRs, BRs, and PFRs. Conventional physics-informed learning begins with a robust foundation but stalls in its learning trajectory due to the limited additional information from the constrained number of shots, eventually matching normal data-driven learning performance with approximately 25 shots. More importantly, consistent with earlier findings, it is more challenging to train BRs using PI-based methods compared to CSTRs. Reptile with normal adaptation requires approximately 40 shots to achieve satisfactory performance (i.e., reaching an MSE magnitude of $10^{-3}$). Conversely, Reptile with physics-informed adaptation attains a similar level with only about 25 shots. Notably, Reptile with physics-informed adaptation consistently maintains the lowest MSE as the number of shots increases.
@@ 5.3 Few-Shot Adaptation to Unseen PFRs
5.4 Exemplary trajectory plots
To further validate the practical significance of the MSE results, we have included exemplary trajectory plots following a 10-shot adaptation to an unseen CSTR, BR, and PFR, as illustrated in Fig. 9. These trajectory plots align with the MSE results, demonstrating that our Reptile approach with physics-informed adaptation achieves the fastest adaptation to the unseen reactor.
(a) 10-shot concentration trajectory plots of an unseen CSTR given a random initial state.
(b) 10-shot temperature trajectory plots of an unseen CSTR given a random initial state.
(c) 10-shot concentration trajectory plots of an unseen BR given a random initial state.
(d) 10-shot temperature trajectory plots of an unseen BR given a random initial state.
(e) 10-shot concentration trajectory plots of an unseen PFR given a random initial state.
(f) 10-shot temperature trajectory plots of an unseen PFR given a random initial state.
Figure 9: Exemplary trajectory plots after 10-shot adaptation to an unseen CSTR, BR, and PFR.
@@ 5.4 Exemplary trajectory plots
Comparing Fig. 6(b), Fig. 7(b), and Fig. 8(b), we can observe that the model adapts more efficiently to certain reactor types over others during training and fine-tuning. Specifically, our observations indicate that the “foundation model” adapts more quickly to CSTRs, followed by BRs, and then PFRs, suggesting that it learns representations of CSTRs more effectively. As a result, during the adaptation phase, CSTRs achieve the best performance, followed by PFRs, and then BRs. The performance of the different methods is summarized as follows: Reptile with physics-informed adaptation $>$ Reptile with normal adaptation $>$ Physics-informed learning $>$ Normal data-driven learning $>$ Transfer learning.
@@ 6 Extension to Varying Integer Orders of Reactions
Next, we extend the proposed method to handle varying integer reaction orders. Specifically, we meta-train “foundation models” for different integer orders of reactions. Consequently, for an unseen integer order of reaction, we integrate their capabilities through ensemble learning, referred to as the ensemble model, with just a few data samples, without requiring precise knowledge of the reaction order, the reaction type, or the complete material and energy balance equations. The high-level system architecture of the ensemble model for varying integer reaction orders is illustrated in Fig. 10. It can be described in three steps: (1) Train $N$ separate “foundation model”s, each corresponding to a different integer reaction order; (2) For a new reaction with an integer reaction order, perform few-shot adaptation on all $N$ “foundation models” to obtain $N$ intermediate models; (3) Apply min-voting (i.e., select the model with the lowest validation/testing MSE) to determine the final ensemble model, utilizing ensemble learning.
@@
    Incorrect order Reptile with normal adaptation: In this approach, after obtaining the “foundation model” via Reptile, an incorrect order of the “foundation model” (i.e., the reaction order used during meta-training differs from that of the unseen meta-testing data) is applied during the meta-testing phase using a standard data-driven adaptation method.
@@
    Incorrect order Reptile with physics-informed adaptation: Similar to the previous approach, an incorrect order of the “foundation model” is used during the meta-testing phase, but this time adaptation is performed using a physics-informed loss function.
@@
Notably, for CSTRs, the ensemble Reptile with normal adaptation performs worse than incorrect-order Reptile with physics-informed adaptation. This is likely because the “foundation model” is better suited to CSTRs than to PFRs and BRs, as observed in Fig. 6(b), Fig. 7(b), and Fig. 8(b). Thus, even with an incorrect reaction order in the “foundation model”, it can still quickly adapt to unseen CSTR reactions. In summary, for a well-trained “foundation model” (particularly with respect to CSTRs), incorporating a physics-informed loss term accelerates adaptation. However, for a “foundation model” that is not as well trained (e.g., with respect to BRs and PFRs), the physics-informed loss term can potentially decelerate adaptation, leading to increased confusion for the model.
@@
To further extend the proposed “foundation model” to incorporate varying fractional reaction orders, the meta-training dataset can be expanded to include random fractional reaction orders. During the meta-testing phase, physics-informed adaptation is then applied to facilitate accurate learning and adaptation.
@@ 7 Limitations and Future Works
Our study demonstrates the model’s ability to adapt to unseen reaction kinetics within the studied reactor types - CSTRs, BRs, and PFRs. However, we do not claim universal applicability to all chemical processes in this work. Achieving full generalizability would require pretraining on a significantly broader set of reaction systems and reactor configurations. This work serves as an initial framework toward that goal, with future research aimed at expanding the training dataset to encompass a more diverse range of chemical processes.
@@
The efficacy of our proposed method is heavily influenced by the challenges associated with training PI-based methods for specific tasks due to the use of estimated parameter values in the physics-informed loss terms. Notably, BRs and PFRs pose greater difficulty for PI-based training, resulting in less discernible differences between our Reptile-based “foundation model” with physics-informed adaptation and the one with normal data-driven adaptation, compared to CSTRs, where the distinction is more pronounced. To address this, inspired by numerical integration, we propose integrating the driving equations directly into the recurrent cells as a different way to incorporate physical prior into the model, rather than incorporating them into the loss term as a soft constraint. Another challenge with PI-based methods, as discussed in Remark 14, lies in the selection of collocation points. Furthermore, our focus on the first discretization element on PFR was primarily driven by ease of validation and computational efficiency. While analyzing behavior further along the reactor could yield additional insights and enrich the dataset, it is beyond the scope of this study (the focus of this work is not to explore methods for data generation or simulate various reactor configurations under different conditions).
Moreover, the current reactor model is idealized and assumes single-phase homogeneous mixing with a simple $A\rightarrow B$ reaction. Several limitations should be noted. The model does not account for multi-step reactions, parallel pathways, or recycling streams, which are common in industrial settings. The hydrodynamics are also simplified, as PFRs are modeled with unidirectional flow, without considering backflow or axial dispersion. Furthermore, the reaction kinetics are assumed to be intrinsic and ideal, without accounting for mass transfer limitations. Another key limitation is the lack of explicit temperature dependency, meaning that variations in density, enthalpy, and reaction kinetics due to temperature fluctuations are not directly modeled. Additionally, this study is simulation-based due to the lack of publicly available industrial datasets. Unlike fields such as computer vision or natural language processing, where large-scale datasets are widely shared, chemical engineering data is often proprietary, making large-scale data collection and model training more challenging.
Future work will focus on improving the model’s accuracy and applicability by addressing these limitations. Expanding to more complex reactor systems (see Fig. 12), such as catalytic fixed-bed reactors and fluidized bed reactors, is an important direction, particularly in systems with multiphase flow, catalyst deactivation, and strong transport-reaction coupling. Incorporating hybrid modeling approaches and physics-based constraints may further enhance generalization to these more intricate reactor types. Additionally, extending the model to handle more realistic reaction networks, including parallel and series reactions, competing pathways, and interacting species, will be crucial for broader applicability. Another avenue for improvement is the explicit modeling of temperature-dependent properties by integrating empirical correlations or physics-informed constraints to capture variations in key physical and chemical parameters.
To further improve generalizability, future research will also leverage Aspen HYSYS, an industry-grade process simulator, to generate additional training data from high-fidelity chemical process models. Furthermore, validating the model on real-world experimental data will be critical in assessing its adaptability to previously unseen reactor types with unknown reaction orders. Finally, while minimizing MSE is a key performance metric, real-world applications must account for noise tolerance. Many downstream tasks, such as neural network-based model predictive control, are inherently robust to measurement noise, meaning that small variations in MSE do not necessarily translate to significant performance differences. Future studies will further evaluate the trade-offs between prediction accuracy and practical usability in noisy environments. By addressing these challenges, we aim to enhance the “foundation model’s” adaptability, making it more applicable to real-world chemical processes beyond the reactor systems studied in this work.
Figure 12: High-level system architecture of future “foundation model” for various chemical reactors.
@@ 8 Conclusion
In this study, we take a step toward developing foundation models for nonlinear chemical process modeling in reactors. By leveraging meta-learning and physics-informed adaptation, we propose a framework that enables rapid adaptation to previously unseen chemical reactions in CSTRs, BRs, and PFRs with minimal training data. Our approach demonstrates improved few-shot adaptation compared to conventional methods, including data-driven learning, physics-informed learning, transfer learning, and standard meta-learning. While this work does not claim to fully establish a foundation model, it highlights a promising direction for improving generalizability in chemical reactor modeling. This framework has the potential to reduce data requirements and enhance adaptability, which could benefit applications such as process optimization and control in chemical engineering.
@@ 9 Acknowledgments
Financial support from the MOE-Tier 1 Grant (A-8002899-00-00) is gratefully acknowledged.